OpenStudy (anonymous):

Lori Bray drives a parcel delivery truck. Her usual route consists of 10 stops. How many ways can Lori travel to these stops and return to her warehouse? A. 90 B. 3,628,800 C. 362,880 D. 10

OpenStudy (anonymous):

\[10!=10\times 9\times 8\times 7\times 6\times 5\times 4\times 3\times 2\times 1\]

OpenStudy (anonymous):

thanx